OpenStudy (anonymous):

Find the slope of the line that passes through the points (3, 6) and (5, 3). -3/2 3/2 2/3

OpenStudy (anonymous):

slope formula (y2-y1)/(x2-x1) (3, 6) and (5, 3). (x1,y1) (x2,y2) plug them into the formula and solve

OpenStudy (anonymous):

I have 3/2

OpenStudy (anonymous):

that is right

OpenStudy (anonymous):

it should be negative

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

they are both negatives so they cancel out Ms_samosino

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so it would be a positive answer

OpenStudy (anonymous):

-3/-2 = 3/2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the slope of the line that passes through the points (-2, 4) and (-5, -6). -2/7 10/3 is my answer, am I correct? -2/3

OpenStudy (anonymous):

(y2-y1)/(x2-x1) (3, 6) and (5, 3). (x1,y1) (x2,y2) (3-6)/(5-3) = -3/2 is it 5-3= -2 it should be positive 2